OpenStudy (anonymous):

You want to build a right triangular garden on the side of your house. The hypotenuse of the triangle measures 8 feet. The other two sides are equal. What are the lengths of the two sides?

OpenStudy (campbell_st):

using pythagoras' theorem  x^2 + x^2 = 8^2 2x^2 = 64 x^2 = 32 \[x = \sqrt{32} = 4\sqrt{2}\]

OpenStudy (anonymous):

thank you :D

OpenStudy (anonymous):

thank you